# Learning to Learn for Few-shot Continual Active Learning

### v3

@@ Abstract
Continual learning strives to ensure stability in solving previously seen tasks while demonstrating plasticity in a novel domain. Recent advances in CL are mostly confined to a supervised learning setting, especially in NLP domain. In this work, we consider a few-shot continual active learning (CAL) setting where labeled data are inadequate, and unlabeled data are abundant but with a limited annotation budget. We propose a simple but efficient method, called Meta-Continual Active Learning. This method sequentially selects the most informative examples from a pool of unlabeled data and requests labels to enhance performance. Specifically, we employ meta-learning and experience replay to address inter-task confusion and catastrophic forgetting. We further incorporate textual augmentations to ensure generalization. We conduct extensive experiments on benchmark text classification datasets to validate the effectiveness of the proposed method and analyze the effect of different active learning strategies in few-shot CAL setting. Our experimental results demonstrate that random sampling is the best default strategy for active learning and memory sample selection to solve few-shot CAL problems.
Index Terms:
continual learning, meta-learning, few-shot learning, active learning, text classification
I Introduction
@@
Continual learning (CL) aims to address the stability-plasticity dilemma in the context of sequential learning. Stability refers to the ability to alleviate the decline in model performance on previously learned tasks, in other words, preventing catastrophic forgetting [1]. Notably, inter-task confusion [2] is one of the major reasons that cause catastrophic forgetting. It is a phenomenon where the model confuses classes from different tasks. Plasticity refers to the ability to adapt to a new task. However, the majority of existing CL methods presume the availability of labeled data in abundance, deeming it adequate for learning every task. Their performance heavily relies on the large quantity and high quality of labeled data, which is impractical.
@@
In real-world scenarios, labeled data are scarce and unlabeled data are abundant. The cost of annotation, particularly concerning NLP tasks, tends to be prohibitively high. To narrow this gap, we contemplate a setting for continual learning that aligns more closely with real-world scenarios. Herein, we provide a small amount of labeled data and a large amount of unlabeled data with limited annotation budget for each continual learning task. We formulate this problem as a few-shot Continual Active Learning (CAL) process. Concretely, model sequentially selects the most worthwhile data from a pool of unlabeled data and requests labels to enhance performance and solve continual learning problems.
@@
Given the success of meta-learning on solving low plasticity, certain studies [3, 4] extend meta-learning to CL setting. In this work, we introduce a simple but efficient method to solve few-shot CAL problems. Specifically, we employ Model-Agnostic Meta-Learning (MAML) framework with memory replay, casting meta-objective as experience replay. In such a way, model learns an optimal initialization that can fast adapt to all seen tasks, thereby enabling fast adaptation while preventing catastrophic forgetting. In addition, we apply consistency regularization via textual augmentations to perform effective learning in the representation space and ensure both intra- and inter-task generalization.
@@
We conduct extensive experiments on CL benchmark datasets from [5], popularized by [6] in lifelong language learning. This collection of datasets includes five text classification datasets from four diverse domains. We validate the effectiveness of our method with different active learning strategies in a 5-shot CAL setup.
The main contributions of this work are threefold:
  • Our work tackles few-shot continual active learning challenges in the NLP domain. Since the data stream is arbitrary and inherently unpredictable, the label space of a forthcoming task may or may not overlap with that of previous tasks. Hence, we formulate a task-free few-shot continual active learning setting to consider both class-incremental learning (i.e., disjoint label space) and domain-incremental learning (i.e., overlapped label space) in one data sequence.
  • We exploit the advantages of MAML and propose a novel method, namely Meta-Continual Active Learning (Meta-CAL). We deliberately devise meta-objective to align examples from different tasks in current embedding space and subsequently solve problem instances. Especially, we use augmented samples selected from the past as query samples to ensure generalization. As such, the proposed method effectively addresses the problems of inter-task confusion and catastrophic forgetting.
@@
  • In our experiments on various text datasets [5], the proposed method achieves an accuracy of more than 62% while utilizing only 1.6% of seen samples and annotation budgets as low as 500 samples per task. Our study shows that random sampling is the best default strategy for active learning and memory sample selection in the few-shot CAL setting. We observe that random sampling ensures generalization, thereby addressing stability-plasticity dilemma.
The remainder of this paper is organized as follows. We introduce related work in Section II. Preliminaries are described in Section III. Section IV and Section V outline the proposed method. Section VI presents the experimental setup and experimental results. We also include further analysis in Section VII. Finally, Section VIII summarizes this paper and shows the future works.
II Related Work
II-A Learning to learn for CL
@@
II-B Continual Active Learning
Continual active learning aims to sequentially labels informative data to maximise model performance and solve continual learning problems. It defines a continual learning problem where the available labeled data are insufficient and the annotation budgets are limited. CASA [14] detects new pseudo-domains and selected data from new pseudo-domains for annotation, while it revisits labeled samples to address catastrophic forgetting. [15] proposes a method to address few-shot CAL (FoCAL). They use a Guaussian mixture model (GMM) for active learning and pseudo-rehearsal for CL, bypassing the need to store real past data. However, neither of these methods addresses continual active learning in NLP. CAL-SD [16] tackles NLP tasks, and uses model distillation to augment memory replay with a diversity and uncertainty AL strategies. To date, continual active learning remains understudied, especially in NLP.
III Preliminaries
In this work, we focus on an online task-free class-incremental learning scenario, in which training data stream passes only once without the presence of “task boundaries” [17]. Based on the task-free setting, we formulate the problem of few-shot continual active learning as follows. Assume that training stream consists of $T$ tasks, $\{\mathcal{T}_{1},\mathcal{T}_{2},...,\mathcal{T}_{t},...,\mathcal{T}_{T}\}$. Each task $\mathcal{T}_{t}$ has a small set of labeled instances $\mathcal{D}^{label}_{t}=\{(x_{i},y_{i})\}_{i=1}^{N_{t}\times K}$ and a pool of unlabeled data $\mathcal{D}^{pool}_{t}=\{(u_{i})\}_{i=1}^{|\mathcal{D}^{pool}_{t}|}$. The labeled set consists of $N_{t}$ classes, each with $K$ samples.
III-A Label Space
@@
III-B Annotation Constraint
We consider each task $\mathcal{T}_{t}$ to have an equal annotation budget $B_{A}$. An acquisition function $a(\cdot)$ dynamically selects informative data points for annotation until acquisition process reaches annotation budget. We denote the selected samples for annotation as $a(\mathcal{D}^{pool}_{t},B_{A})$ and newly annotated sample set as $\mathcal{D}^{new}_{t}$.
III-C Memory Constraint
Replay-based model revisits a small subset of labeled data $\mathcal{D}_{1:1-t}$ to regularize model $f$ while learning $\mathcal{T}_{t}$. We limit the amount of past training data saved in memory $\mathcal{M}$, which should not exceed memory budget threshold $B_{\mathcal{M}}$.
III-D Objectives
Assume we have a learner $f_{\theta}$ and current task $\mathcal{T}_{t}$, our learning objectives are: (a) perform adaptation on $\mathcal{D}_{t}=\mathcal{D}^{label}_{t}\bigcup\mathcal{D}^{new}_{t}$, i.e., plasticity:
@@ III-D Objectives
where $\mathcal{L}_{\mathcal{T}_{t}}$ is the task loss, $\theta_{t}$ is the initial state and $\theta_{t}=\tilde{\theta}_{t-1}$; (b) prevent inter-task confusion and catastrophic forgetting, i.e., stability:
@@
III-E Few-shot CAL Challenges
The few-shot CAL setting poses hurdles for the optimization of neural networks. The main challenges are:
  • Quickly adapting the model to the current task while alleviating catastrophic forgetting, with limited labeled examples;
  • Addressing inter-task confusion with data scarcity;
  • Exploiting an efficient active learning strategy that makes the best use of available data to solve continual learning problems.
III-F Active Learning Strategies
@@ IV Learning to Learn for CAL
IV-A Learning to Fast Adapt
We approximately align the meta-objective with the objectives shown in Eqn. 1 and Eqn. 2 as
@@ IV-A Learning to Fast Adapt
[…]
where $U_{\mathcal{T}_{t},\mathcal{D}_{t}}(\theta_{t})$ is update operation on $\theta_{t}$ using training set $\mathcal{D}_{t}$. Specifically, $U_{\mathcal{T}_{t},\mathcal{D}_{t}}(\theta_{t})$ describes SGD optimization acting on $\theta_{t}$ as
@@
Hereby, instead of finding finetuned $\tilde{\theta}_{t}$, model $f$ learns an optimal initialization $\theta_{t}$ that effectively adapts to $\mathcal{D}_{1:t}$ with few labeled examples.
IV-B Learning to Continually Learn
In general, the full datasets $\mathcal{D}_{1:t-1}$ are not available while learning $\mathcal{D}_{t}$. Thus, we retrieve past samples from memory set $\mathcal{M}$ for experience replay in meta-objective,
@@ IV-B Learning to Continually Learn
[…]
In such a way, we leverage the selected examples from the past to constrain the learning behaviour of $f$. As a result, it effectively tackles the problems of catastrophic forgetting and inter-task confusion.
IV-C Learning to Generalize
We also exploit data augmentations to enhance model’s ability to generalize. Especially, we apply consistency regularization [25], which is built on the assumption that perturbations of the same input should not affect the output. Inspired by FixMatch [26], we employ two types of data augmentations in meta-training, strong and weak, denoted by $\mathcal{A}(\cdot)$ and $\alpha(\cdot)$, respectively. In contrast to FixMatch, we employ textual augmentations under full supervision to ensure generalization with limited data availability. Specifically, weak and/or strong augmentations are applied in the inner loop to enhance intra-task generalization, while strong augmentations are used in the outer loop to improve both intra- and inter-task generalization. More details are shown in §V-B.
V Model
V-A Model Architecture
@@ V-B Training
Inner-loop optimization
Inner-loop algorithm performs task-specific learning. We introduce data augmentations as the regularization term to improve intra-task generalization. The inner loop loss for training samples $\mathcal{D}_{i}^{label}$ at time step $i$ is
$\small\begin{split}\mathcal{L}_{\mathbf{inner}}^{D_{i}^{label}}(\theta)&=\sum_{(x,y)\in D_{i}^{label}}[w\mathcal{L}_{CE}\big{(}f_{\theta}(x),y\big{)}\\ &+(1-w)\mathcal{L}_{CE}\big{(}f_{\theta}(\alpha(x)),y\big{)}]\\ \end{split}$ (8)
where $w$ denotes the relative weight and $\mathcal{L}_{CE}$ is the cross-entropy loss. [Annotation process] When the received training batches are unlabeled, we apply acquisition function $a(\mathcal{D}_{i}^{pool},m\cdot n_{a})$ to select informative data points for annotation. The selection size per batch for annotation $n_{a}=\lceil\frac{b\cdot B_{A}}{|\mathcal{D}^{pool}|}\rceil$ where $b$ is the batch size and $B_{A}$ is the annotation budget. Then, the inner-loop loss for newly annotated training batches $\mathcal{D}^{new}_{i}$ is
$\begin{split}\mathcal{L}_{\mathbf{inner}}^{\mathcal{D}^{new}_{i}}(\theta)&=\sum_{(x,y)\in\mathcal{D}^{new}_{i}}[w\mathcal{L}_{CE}\big{(}f_{\theta}(x),y\big{)}\\ &+(1-w)\mathcal{L}_{CE}\big{(}f_{\theta}(\mathcal{A}(x)),f_{\theta}(\alpha(x))\big{)}]\\ \end{split}$ (9)
@@
[Memory sample selection] We dynamically update memory buffer $\mathcal{M}$ using reservoir sampling to ensure generalization while avoiding overfitting. Reservoir sampling [3] randomly selects a fixed number of training samples without knowing the total number of samples in advance. We use reservoir sampling to select $n_{s}$ examples per class on the incoming data stream $\mathcal{D}^{label}_{i}\bigcup\mathcal{D}^{new}_{i}$ with an equal selection probability for all data seen so far. Note that the current label space $\mathcal{Y}_{i}$ might overlap with previous label spaces, $\mathcal{Y}_{i}\cap(\mathcal{Y}_{1}\cup\mathcal{Y}_{2}\cup\ldots\cup\mathcal{Y}_{i-1})\neq\emptyset$. We automatically update memory samples for $\mathcal{Y}_{i}$ to maintain a fixed amount of memory samples per class.
Input: Initial parameters $\theta=\theta_{\mathrm{e}}\cup\theta_{\mathrm{clf}}$, training set $\mathcal{D}^{train}=\mathcal{D}^{label}\cup\mathcal{D}^{pool}$, memory buffer $\mathcal{M}$, acquisition function $a(\cdot)$, inner-loop learning rate $\alpha$, outer-loop learning rate $\beta$, and memory samples per class $n_{s}$ and annotation budget per task is $B_{A}$.
Output: Trained initial parameters $\theta$ and Memory $\mathcal{M}$
1 for $i=1,2$, … do
2       Receiving $m$ batches of examples, $\mathcal{D}_{i}^{train}$, from the stream
3       if $\mathcal{D}_{i}^{train}$ from a unseen task then
4             Compute selection size per batch for annotation as $n_{a}=\lceil\frac{b\cdot B_{A}}{|\mathcal{D}^{pool}|}\rceil$
5       end if
6      [Inner Loop]
7       // Trained on Labeled data first for each task
8       if $\mathcal{D}_{i}^{train}\subseteq\mathcal{D}^{label}$  then
9             $\mathcal{D}_{i}^{label}\leftarrow\mathcal{D}_{i}^{train}$ and apply weak augmentation on samples from $\mathcal{D}_{i}^{label}$
10             Perform SGD on $\theta_{\mathrm{clf}}$ to minimize Eqn.8
11       else if $\mathcal{D}_{i}^{train}\subseteq\mathcal{D}^{pool}$ then
12             $\mathcal{D}_{i}^{pool}\leftarrow\mathcal{D}_{i}^{train}$ and do annotation per batch as $\mathcal{D}_{i}^{new}\xleftarrow{label}a(\mathcal{D}_{i}^{pool},m\cdot n_{a})$
13             Apply both weak and strong augmentations on samples from $\mathcal{D}_{i}^{new}$
14             Perform SGD on $\theta_{\mathrm{clf}}$ to minimize Eqn.9
15       end if
16      Select memory sample using ReservoirSampling$(\mathcal{D}^{label}_{i}\bigcup\mathcal{D}^{new}_{i},\mathcal{M})$ and update $\mathcal{M}$
17       [Outer Loop]
18       Read ALL examples from $\mathcal{M}$ and apply strong augmentation
19       Perform Adam update on $\theta$ to minimize Eqn.12
20       if all training data are seen then
21             Stop training
22       end if
23
24 end for
@@ 21             Stop training
TABLE I: Average accuracy of four training set orders for sequentially learning five tasks from four domains
Method Save Ratio AL AVG.

### v4

@@ Abstract
Continual learning strives to ensure stability in solving previously seen tasks while demonstrating plasticity in a novel domain. Recent advances in continual learning are mostly confined to a supervised learning setting, especially in NLP domain. In this work, we consider a few-shot continual active learning setting where labeled data are inadequate, and unlabeled data are abundant but with a limited annotation budget. We exploit meta-learning and propose a method, called Meta-Continual Active Learning. This method sequentially queries the most informative examples from a pool of unlabeled data for annotation to enhance task-specific performance and tackle continual learning problems through meta-objective. Specifically, we employ meta-learning and experience replay to address inter-task confusion and catastrophic forgetting. We further incorporate textual augmentations to avoid memory over-fitting caused by experience replay and sample queries, thereby ensuring generalization. We conduct extensive experiments on benchmark text classification datasets from diverse domains to validate the feasibility and effectiveness of meta-continual active learning. We also analyze the impact of different active learning strategies on various meta continual learning models. The experimental results demonstrate that introducing randomness into sample selection is the best default strategy for maintaining generalization in meta-continual learning framework.
keywords:
continual learning, meta-learning, active learning, few-shot learning, text classification
1 Introduction
@@
Continual Learning (CL) aims to address the stability-plasticity dilemma in the context of sequential learning. Stability refers to the ability to alleviate the decline in model performance on previously learned tasks, in other words, preventing catastrophic forgetting [22]. Notably, inter-task confusion [15] is one of the major reasons that causes catastrophic forgetting. It is a phenomenon where the model confuses classes from different tasks. Plasticity refers to the ability to adapt to a new task. However, the majority of existing CL methods presume the availability of labeled data in abundance, deeming it adequate for learning every task. Their performance heavily relies on the large quantity and high quality of labeled data, which is impractical.
@@
In real-world scenarios, labeled data are scarce while unlabeled data are abundant. The cost of annotation, particularly in the field of Natural Language Processing (NLP), tends to be prohibitively high. To narrow this gap, we address the problems in a setting that more closely aligns with real-world scenarios, namely few-shot Continual Active Learning (CAL) [2]. In this setting, only a small subset of labeled data is provided for each task with a limited annotation budget. In such a way, model should sequentially select the most worthwhile examples from a pool of unlabeled data and requests labels to enhance performance and simultaneously solve continual learning problems. Introducing active learning into the scheme can be challenging as active learning techniques are typically designed to query from a static data distribution. They may not be able to dynamically capture most relevant samples for queries to prevent catastrophic forgetting.
Replay-based methods have been shown to be particularly effective for NLP tasks [37]. These methods often retain a certain amount of past samples to prevent catastrophic forgetting. Consequently, they are prone to memory over-fitting, especially when the availability of labeled examples is limited. However, active learning can further escalate the problem by biasing the sample selection towards the memory set. Hence, integrating active learning techniques into CL model can be quite challenging.
@@
Given the success of meta-learning on solving low plasticity, certain studies [29, 14] extend meta-learning to CL setting. In this work, we exploit the advantages of meta-learning and uses Model-Agnostic Meta-Learning (MAML) [13] framework to solve few-shot CAL problems, called Meta-Continual Active Learning. By applying active learning for task-specific tuning and casting meta-objective as experience replay, the learning objective is deliberately formulated to learn an optimal or a suboptimal initial model state that can rapidly adapt to a balanced subset of all encountered task. Thereby, this method enables fast adaptation while preventing catastrophic forgetting in few-shot learning. In addition, we apply consistency regularization via textual augmentations to address memory overfitting problems that inherent in replay-based method and exacerbated by active learning acquisition.
@@
We conduct extensive experiments on benchmark datasets from [41], popularized by [12] in lifelong language learning. This collection of datasets includes five text classification datasets from four diverse domains. We demonstrate the effectiveness of the proposed framework in a 5-shot CAL setup. This paper also examines how various active learning approaches impact the performance of meta continual learning models.
The main contributions of this paper are fourfold:
  • Leveraging the strengths of meta-learning, we introduce an optimization-based method, namely Meta-Continual Active Learning (Meta-CAL). This method reformulates the meta-objective such that it learns an optimal or a suboptimal initial model state that can effectively adapt to all seen tasks. Thereby, it provides a solution to inter-task confusion and catastrophic forgetting even with limited availability of labeled samples.
  • We integrate active learning into the proposed framework to enhance task-specific tuning. This allows the model to dynamically and selectively query the most informative samples from a pool of unlabeled data, thereby improving performance in a resource-constrained scenario.
  • To address inevitable memory overfitting problems caused by experience replay and active learning, we apply consistency regularization to meta examples via data augmentations. This further ensures intra- and inter-task generalization.
@@
  • In the experiments, the proposed method achieves an accuracy of more than 62% while utilizing only 1.6% of the past samples and maintaining annotation budgets as low as 500 samples for each task. The results demonstrate the feasibility and effectiveness of meta-continual active learning. Furthermore, we observe that random sampling facilitates generalization in meta continual learning, thereby addressing the stability-plasticity dilemma.
2 Related Work
2.1 Continual Learning
Existing approaches can be categorised into three main mainstreams, i.e., regularization-based methods [20, 21], replay-based methods [12, 17] and architecture-based methods [3, 40].
Regularization-based methods often impose a penalty or regularization term to the loss function. These methods often penalise changes in non-trivial parameters or constrain the variations in gradients learned from previous tasks. However, most methods for NLP show a preference for replay-based approaches [37] to avoid unexpected output from tuning the parameters of deep neural networks [38].
Replay-based methods, also known as rehearsal-based methods or memory-based methods, involves revisiting a small amount of past samples (i.e., experience replay) or generating pseudo past samples while adapting to new domain. The popular retrieval schemes for experience replay are random [8, 29, 12, 18], K-Means [38, 16] and Mean-of-Feature [28, 9].
Architecture-based methods dynamically change the model architecture to learn a new task. In general, these methods preserve or partially preserve past fine-tuned parameters and introduce task-specific parameters for the new domain. However, it is challenging to manage the scale of the model since the parameter size continuously accumulates with the increase in the number of seen tasks.
2.2 Meta Continual Learning
@@
2.3 Continual Active Learning
Continual active learning aims to sequentially labels informative data to maximise model performance and solve continual learning problems. It defines a continual learning problem where the available labeled data are insufficient and the annotation budgets are limited. CASA [27] detects new pseudo-domains and selected data from new pseudo-domains for annotation, while it revisits labeled samples to address catastrophic forgetting. [2] propose a method to address few-shot CAL (FoCAL). They use a Guaussian mixture model (GMM) for active learning and pseudo-rehearsal for CL, bypassing the need to store real past data. However, neither of these methods addresses continual active learning in NLP. CAL-SD [10] tackles NLP tasks, and uses model distillation to augment memory replay with a diversity and uncertainty AL strategies. To date, continual active learning remains understudied, especially in NLP.
3 Preliminaries
In this work, we focus on task-free class-incremental learning scenario, where training data stream passes only once without the presence of “task boundaries” [34]. Based on the task-free setting, we formulate the problem of few-shot continual active learning as follows. Assume that training stream consists of $T$ tasks, $\{\mathcal{T}_{1},\mathcal{T}_{2},...,\mathcal{T}_{t},...,\mathcal{T}_{T}\}$. Each task $\mathcal{T}_{t}$ contains $N_{t}$-way $K$-shot set $\mathcal{D}^{label}_{t}=\{(x_{i},y_{i})\}_{i=1}^{N_{t}\times K}$ and a pool of unlabeled data $\mathcal{D}^{pool}_{t}=\{(u_{i})\}_{i=1}^{|\mathcal{D}^{pool}_{t}|}$.
3.1 Label Space
@@
3.2 Annotation Constraint
We consider each task $\mathcal{T}_{t}$ to have an equal annotation budget $B_{A}$. An acquisition function $a(\cdot)$ dynamically query informative data points for annotation until acquisition process reaches annotation budget. We denote the selected samples for annotation as $a(\mathcal{D}^{pool}_{t},B_{A})$ and newly annotated sample set as $\mathcal{D}^{new}_{t}$.
3.3 Memory Constraint
Replay-based model revisits a small subset of labeled data $\mathcal{D}_{1:1-t}$ to regularize model $f$ while learning $\mathcal{T}_{t}$. We limit the amount of past training data saved in memory buffer $\mathcal{M}$, which should not exceed memory budget threshold $B_{\mathcal{M}}$.
3.4 Objectives
Assume we have a learner $f_{\theta}$ and current task $\mathcal{T}_{t}$, the learning objectives are: (a) perform adaptation on $\mathcal{D}_{t}=\mathcal{D}^{label}_{t}\bigcup\mathcal{D}^{new}_{t}$, i.e., plasticity:
@@ 3.4 Objectives
where $\mathcal{L}_{\mathcal{T}_{t}}$ is the task loss, $\theta_{t}$ is the initial state and $\theta_{t}=\tilde{\theta}_{t-1}$; (b) prevent inter-task confusion and catastrophic forgetting of prior tasks, i.e., stability:
@@
3.5 Active Learning Strategies
@@ 4 Learning to Learn for CAL
4.1 Learning to Fast Adapt
We approximately align the meta-objective with the objectives shown in Eqn. 1 and Eqn. 2 as follows,
@@ 4.1 Learning to Fast Adapt
s.t. $\tilde{\theta}_{t}=U_{\mathcal{T}_{t},\mathcal{D}_{t}}(\theta_{t})$
where $U_{\mathcal{T}_{t},\mathcal{D}_{t}}(\theta_{t})$ is update operation on $\theta_{t}$ using training set $\mathcal{D}_{t}$. Specifically, $U_{\mathcal{T}_{t},\mathcal{D}_{t}}(\theta_{t})$ describes optimization acting on $\theta_{t}$ as
@@
Hereby, instead of finding tuned $\tilde{\theta}_{t}$, model $f$ learns an optimal initialization $\theta_{t}$ that can effectively adapt to $\mathcal{D}_{1:t}$ with few labeled examples.
4.2 Learning to Continually Learn
In general, the full datasets $\mathcal{D}_{1:t-1}$ are not available while learning $\mathcal{D}_{t}$. Thus, we retrieve a small subset of past samples $\mathcal{M}$ for experience replay in meta-objective,
@@ 4.2 Learning to Continually Learn
s.t. $\tilde{\theta}_{t}=U_{\mathcal{T}_{t},\mathcal{D}_{t}}(\theta_{t})$
$=\theta_{t}-\alpha\nabla_{\theta_{t}}\mathbb{E}_{(x,y)\sim\mathcal{D}_{t}}\mathcal{L}(f_{\theta_{t}}(x),y)$
In such a way, we leverage the selected examples from the past to constrain the learning behaviour of $f$. As a result, it effectively tackles the problems of inter-task confusion and catastrophic forgetting.
4.3 Learning to Generalize
We also exploit data augmentations to enhance model’s ability to generalize. Especially, we apply consistency regularization [4], which is built on the assumption that perturbations of the same input should not affect the output. Inspired by FixMatch [31], we employ two types of data augmentations in meta-training, strong and weak, denoted by $\mathcal{A}(\cdot)$ and $\alpha(\cdot)$, respectively. In contrast to FixMatch, we employ textual augmentations under full supervision to ensure generalization with limited data availability. Specifically, weak and/or strong augmentations are applied in the inner loop to enhance intra-task generalization, while strong augmentations are used in the outer loop to improve both intra- and inter-task generalization. More details are shown in §5.2.
5 Model
5.1 Model Architecture
@@ 5.2 Training
5.2.1 Inner-Loop Optimization
Inner-loop algorithm performs task-specific tuning. We introduce data augmentations as the regularization term to improve intra-task generalization. The inner loop loss for training samples $\mathcal{D}_{i}^{label}$ at time step $i$ is
$\small\mathcal{L}_{\mathbf{inner}}^{D_{i}^{label}}(\theta)=\sum_{(x,y)\in D_{i}^{label}}[w\mathcal{L}_{CE}(f_{\theta}(x),y)+(1-w)\mathcal{L}_{CE}\big{(}f_{\theta}(\alpha(x)),y\big{)}]$ (6)
where $w$ denotes the relative weight and $\mathcal{L}_{CE}$ is the cross-entropy loss.
Annotation Process
When the received training batches are unlabeled, we apply acquisition function $a(\mathcal{D}_{i}^{pool},m\cdot n_{a})$ to select informative data points for annotation. The selection size per batch for annotation $n_{a}=\lceil\frac{b\cdot B_{A}}{|\mathcal{D}^{pool}|}\rceil$ where $b$ is the batch size and $B_{A}$ is the annotation budget.
Then, the inner-loop loss for newly annotated training batches $\mathcal{D}^{new}_{i}$ is
$\small\mathcal{L}_{\mathbf{inner}}^{\mathcal{D}^{new}_{i}}(\theta)=\sum_{(x,y)\in\mathcal{D}^{new}_{i}}[w\mathcal{L}_{CE}\big{(}f_{\theta}(x),y\big{)}+(1-w)\mathcal{L}_{CE}\big{(}f_{\theta}(\mathcal{A}(x)),f_{\theta}(\alpha(x))\big{)}]$ (7)
@@ 5.2.1 Inner-Loop Optimization
Memory Sample Selection
We dynamically update memory buffer $\mathcal{M}$ using reservoir sampling to ensure generalization while avoiding overfitting. Reservoir sampling [29] randomly selects a fixed number of training samples without knowing the total number of samples in advance. We use reservoir sampling to select $n_{s}$ examples per class on the incoming data stream $\mathcal{D}^{label}_{i}\bigcup\mathcal{D}^{new}_{i}$ with an equal selection probability for all data seen so far. Note that the current label space $\mathcal{Y}_{i}$ might overlap with previous label spaces. We automatically update memory samples for $\mathcal{Y}_{i}$ to maintain a fixed amount of memory samples per class.
@@
1:Training set $\mathcal{D}^{train}=\mathcal{D}^{label}\cup\mathcal{D}^{pool}$, acquisition function $a(\cdot)$, inner-loop learning rate $\alpha$, outer-loop learning rate $\beta$, memory samples per class $n_{s}$, annotation budget per task $B_{A}$
2:Trained model with parameters $\theta=\theta_{\mathrm{e}}\cup\theta_{\mathrm{clf}}$ and memory buffer $\mathcal{M}$
3:for $i=1,2,\ldots$ do
4:    Receiving $m$ batches of examples, $\mathcal{D}_{i}^{train}$, from the stream
5:    if $\mathcal{D}_{i}^{train}$ is from an unseen task then
6:         Compute selection size per batch for annotation as $n_{a}=\lceil\frac{b\cdot B_{A}}{|\mathcal{D}^{pool}|}\rceil$
7:    end if
8:    [Inner Loop] $\triangleright$ Trained on Labeled data first for each task
9:    if $\mathcal{D}_{i}^{train}\subseteq\mathcal{D}^{label}$ then
10:         $\mathcal{D}_{i}^{label}\leftarrow\mathcal{D}_{i}^{train}$ and apply weak augmentation on samples from $\mathcal{D}_{i}^{label}$
11:         Perform SGD on $\theta_{\mathrm{clf}}$ to minimize Eqn. 6
12:    else if $\mathcal{D}_{i}^{train}\subseteq\mathcal{D}^{pool}$ then
[…]
14:         Apply both weak and strong augmentations on samples from $\mathcal{D}_{i}^{new}$
15:         Perform SGD on $\theta_{\mathrm{clf}}$ to minimize Eqn. 7
16:    end if
17:    Select memory sample using ReservoirSampling$(\mathcal{D}^{label}_{i}\bigcup\mathcal{D}^{new}_{i},\mathcal{M})$ and update $\mathcal{M}$
18:    [Outer Loop]
19:    Read ALL examples from $\mathcal{M}$ and apply strong augmentation
20:    Perform Adam update on $\theta$ to minimize Eqn. 10
21:    if all training data are seen then
22:         Stop training
23:    end if
24:end for
5.2.2 Outer-Loop Optimization
Outer-loop algorithm optimizes initial parameter set $\theta$ to a setting that $f$ can effectively adapt to $\mathcal{D}_{1:t}$ via a few gradient updates. Due to memory constraints, we only have a few amount of samples from the past. The model reads all examples from memory buffer $\mathcal{M}$. Then, the outer-loop objective is to have $\tilde{\theta}=\theta_{\mathrm{e}}\cup\tilde{\theta}_{\mathrm{clf}}$ generalize well on $\mathcal{D}_{1:t}$ using $\mathcal{M}$, shown in Eqn. 5,
To reduce the complexity of the second-order computation in outer loop, we use first-order approximation, namely FOMAML. The outer-loop optimization process is
5.3 Testing
where $S\subseteq\mathcal{M}$. And the optimization process is,
6 Experiments
6.1 Datasets
We use the text classification benchmark datasets from [41], including AGNews (news classification; 4 classes), Yelp (sentiment analysis; 5 classes), Amazon (sentiment analysis; 5 classes), DBpedia (Wikipedia article classification; 14 classes) and Yahoo (questions and answers categorization; 10 classes). This collection contains 5 tasks from 4 different domains, which covers class- and domain-incremental learning in task sequence. We randomly sample 5 labeled instances per class, 10,000 unlabeled instances, and 7,600 test examples from each of the datasets. Following prior studies [36, 18, 17], we concatenate training sets in 4 different orderings as shown in Table 1.
Table 1: Input Dataset Orders
Index Dataset Orders
[…]
6.2 Implementation details
6.3 Baselines
Baseline model:
  1. 1.
    MAML-SEQ: Online FOMAML algorithm.
  2. 2.
    OML-ER [18]: OML with 5% episodic experience replay rate [^3] The default 1% rate shows bad performance + reservoir sampling.
  3. 3.
    C-MAML [^4] Being modified for NLP tasks [14]: OML with Meta & CL objective alignment + reservoir sampling.
  4. 4.
    Meta-CAL (ours): OML with Meta & CL objective alignment + consistency regularization + reservoir sampling.
  5. 5.
    FULL: Supervised C-MAML, trained on full datasets.
Memory sample selection:
  1. 1.
    Prototype [17]: Select representative samples that are closest to dynamically updated prototypes in representation space.
  2. 2.
    Ring Buffer [7]: Use a ’First-In, First-Out’ scheme to update buffer.
  3. 3.
    Reservoir Sampling [29]: Randomly select data with an equal selection probability.
6.4 Evaluation Metrics
Based on the prior work [39], we use three comprehensive and widely-used metrics in CL, i.e., accuracy, backward transfer and forward transfer. Let $R_{k,i}$ be the macro-averaged accuracy evaluated on the test set of the $i$ task after sequentially learning $t$ tasks,
Accuracy:
$\mathrm{ACC}_{t}=\frac{1}{t}\sum^{t}_{i=1}R_{t,i}$
Overall accuracy is the weighted average accuracy of all seen tasks $\mathcal{T}_{1:T}$.
Backward transfer (stability evaluation):
$\mathrm{BWT}_{k}=\frac{1}{k-1}\sum^{k-1}_{i=1}(R_{k,i}-R_{i,i})$
BWT measures how the updated parameters affect model performance on all previously seen tasks.
Forward transfer (plasticity evaluation):
$\mathrm{FWT}_{k}=\frac{1}{k-1}\sum^{k}_{i=2}(R_{i,i}-R_{0,i})$
FWT quantifies the average impact of all preceding tasks on the current $k$ task.
6.5 Main Results
In Table 2, we compare various models with four AL strategies. These strategies include Random (denoted as RAND), Representative via KMeans (denoted as REP), Diversity via KMeans (denoted as DIV), and Uncertainty via Least-confidence (denoted as UNC). The label budget $B_{A}$ = 2000 examples per task. Each record is the average of three best results from five runs.
Table 2: Average Accuracy of Four Training Set Orders
Method Save Ratio [^1] The ratio of saved samples to the total number of seen samples AL AVG.